Find a general solution of the ODE by using the method of undetermined coefficients.

24 - 5y + 2y = (t + 3)et/2
Answer #1

The desde of the green cam then 2 As e/2 is present on the role = 0, e2+ Gett. Zy - 56+ 2y = (t+13) e12 - 0 Amilliary equatio